OpenStudy (anonymous):

Given the system of the following linear equations: 5x – 3y = 9 5x + 4y = 23 Solve for x and y. Question 5 options: (-2, 3) (-2, -3) (3, 2) (3, -2)

OpenStudy (the1loveguru):

Easy

OpenStudy (anonymous):

Will Fan and Medal :D

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

A is not correct @The1LoveGuru

OpenStudy (anonymous):

Can you explain @tom982 ?

OpenStudy (the1loveguru):

5x - 3y = 9 5x + 4y = 23 okay so first you have to "eliminate" one of the variables so you can easily "eliminate" X since they have the same coeffiecents BUT they have to have the opposite signs soooo you end up with 5x - 3y = 9 -5x -4y = -23 ( apply the - sign to everything on both sides) then you can get rid of the x and solve for the y so you end up with -7y = -14 y = -14/-7 = +2 THEN you plug back the 2 into the y variable (any ORIGINAL equation is fine) so you end up solving for X 5x -3(2) = 9 5x - 6 = 9 5x = 9 +6 5x = 15 x = 15/5 = +3 (3,2 ) I meant C sorry!

OpenStudy (the1loveguru):

I meant C Idek why I chose A. SOrry.

OpenStudy (anonymous):

\[5x – 3y = 9\]\[5x + 4y = 23\]Rearrange these to have an equal term:\[5x=9+3y\]\[5x=23-4y\]Equate them:\[9+3y=23-4y\]Simplify\[7y=14\]\[y=2\]Use one of the initial equations and our value for \(y\) to find \(x\):\[5x-3y=9\]\[5x-3(2)=9\]\[5x=15\]\[x=3\] Giving our solution: \((3,2)\)

OpenStudy (anonymous):

That makes a lot of sense and it's OK i haven't submitted it yet :)

OpenStudy (the1loveguru):

I am so sorry dude.

OpenStudy (the1loveguru):

I feel so bad right now.